# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 8.2 Parton-level uncertainties
Figure 5: Variations of the simulation of the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ simulation with the nuisance parameters $\nu_{R}$ and $\nu_{F}$, modifying the renormalization and factorization scales, respectively. The dashed lines show the training data, and the solid lines show the prediction from $\hat{S}_{\textrm{scale}}(\boldsymbol{x}|\nu_{R},\nu_{F})$ as obtained from the parametric regression tree. All shapes are normalized to the simulation at the SM parameter point, corresponding to $\nu_{R}=\nu_{F}=0$.
Among the largest systematic effects are the uncertainties in the factorization and renormalization scales, whose modeling is described in detail in Sec. 4.2. According to Eq. 16, values $\nu_{R}=\pm 1$ and $\nu_{F}=\pm 1$ for the nuisance parameters vary the scales $\mu_{R}$ and $\mu_{F}$ by a factor of 2. From the simulation, we obtain event weights for all eight combinations of the scale variations, concretely,
@@
The nominal SM simulation corresponds to $\nu_{R}=\nu_{F}=0$. We model the scale uncertainties up to quadratic accuracy in the nuisance parameter dependence with the ansatz
@@
The five independent functions are labeled by $A={R,F,RR,FF,RF}$ for the two linear, the two quadratic, and a single mixed term. Therefore, the eight variations in $\mathcal{V}$ overconstrain these five functions.
We fit the parametric regression tree according to Sec. 7.1 in the standard configuration described in Sec. 8.1. The result is shown in Fig. 5 in one-dimensional projections. We show the observables with the strongest variations with $\nu_{R}$ and $\nu_{F}$. Correlated scale variations of a factor 2 of $\mu_{R}$ and $\mu_{F}$ correspond to $\nu_{R}=\nu_{F}=\pm 1$ and can reach 8-10%. The exception is the observable $p_{\textrm{T}}$ where the variation exceeds 20% in the tail. In this regime, the fit shows a small deficit of 1-2% compared to the true variations. We conjecture this to be a small residual inflexibility in the quadratic model. Because the parameters $\mu_{R}$ and $\mu_{F}$ only cover uncertainties related to the insufficient perturbative control, we ignore the slight mismatch for the present purpose. All other kinematic features show a smaller dependence on the shape.
@@
We also consider event simulation at the SM parameter point with the alternative powheg generator, providing a similarly sized event sample of the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ process at NLO accuracy in the strong coupling constant. As described in Sec. 5.3, and despite the caveats mentioned there, we associate a nuisance parameter $\nu_{\textrm{POW}}$ with the alternatives where $\nu_{\textrm{POW}}=0$ corresponds to simulation with MadGraph5_amc@nlo and $\nu_{\textrm{POW}}=1$ corresponds to simulation with powheg. Thus, we have $\mathcal{V}=\{1\}$ which allows us to train a single-parameter linear surrogate for the (log-) DCR, denoted by $\hat{S}_{\textrm{POW}}(\boldsymbol{x}|\nu_{\textrm{POW}})$. We show the features’ one-dimensional projections exhibiting shape differences in Fig. 6. Because we train with stochastically independent samples, mild statistical fluctuations are visible in the tails of the training data, which the parametric regression tree averages out.
We leave uncertainties in the PDFs, requiring, e.g., 100 nuisance parameters corresponding to the variations in the PDFs eigendirections [79] to future treatment. Instead, and to account for uncertainties in the efficiency of the $m(\mathrm{t}\overline{\mathrm{t}})$ requirement, we chose a normalization uncertainty of 15%, concretely $\alpha_{\textrm{norm}}=1.15$.
@@ 8.3 Jet energy calibration uncertainties
To assess the effects of the uncertainties in the reconstructed transverse jet momenta, we vary the predicted values from Delphes by the variations provided in Ref. [63]. We use the “total” uncertainty in the CMS JES and propagate the variations to the event selection, the missing energy, the kinematic reconstruction of the top quark kinematics, and the other event features in $\boldsymbol{x}$. The per-jet variations depend on the nominal $p_{\textrm{T}}$ and pseudo-rapidity, which are not part of $\boldsymbol{x}$. Hence, the resulting function $J_{\nu_{\textrm{JES}}}(\boldsymbol{x},\boldsymbol{z})$ is dependent on the latent event configuration. The variations define synthetic data sets according to Sec. 4.4.1, and we chose $\mathcal{V}=\{-1,-0.5,0.5,1\}$ for the nuisance parameter $\nu_{\textrm{JES}}$. It is designed to parameterize the effects of the per-jet variations on $\boldsymbol{x}$ and is given in units of standard deviations of the underlying JES uncertainty. The half-integer values for $\nu_{\textrm{JES}}$ refine the customary choice of ±$\pm 1\sigma$ variations in most binned LHC data analyses. We fit a log-linear surrogate
@@
of the JES dependence. In Fig. 7, we compare the fit of the surrogate to the variations in the training data and find overall excellent agreement. The dependence of the variation on the observable is flat except $p_{\textrm{T}}(\mathrm{t}\overline{\mathrm{t}})$ where it raises from 2%–5%. In the tails of $m(\mathrm{t}\overline{\mathrm{t}})$, the variations in the training data are slightly asymmetric, a feature the linear model can not capture. We observe that the trained surrogate approximately symmetrizes the total uncertainty, and we leave a refinement with a higher-degree surrogate for future treatment.
@@ 8.4 Uncertainties in tagging efficiencies
Uncertainties in the b-tagging efficiency of jets and a recipe for the application are provided in Ref. [63]. It relies on $p_{\textrm{T}}$, the pseudo-rapidity, and a binary nominal b-tag label taken from Delphes. To apply the b-tagging variations, we also need the generator-level flavor $f$ of the jet within the set $\{\textrm{udsg},\mathrm{c},\mathrm{b}\}$. Separately for each flavor, we use the nominal Delphes simulation to parametrize $p_{\textrm{T}}$ and $\eta$-dependent $\mathrm{b}$-tagging efficiencies $\varepsilon_{f}(p_{\textrm{T}},\eta)$. The scale factors $\textrm{SF}_{f}(p_{\textrm{T}},\eta)$ and the variations $\Delta\textrm{SF}_{f}(p_{\textrm{T}},\eta)$ are used to produce variations for two systematic effects. The heavy-flavor (HF) tagging uncertainty covers uncertainties in jets’ tagging rate originating from b or c quarks. For non-zero values of the nuisance $\nu_{\textrm{HF}}$, the efficiencies of the b and c jets are, therefore, changed in a correlated manner. The light-flavor (LF) mistagging uncertainty covers uncertainty in the mistag rate of jets originating from light quarks or gluons. We associate the nuisance parameters $\nu_{\textrm{HF}}$ and $\nu_{\textrm{LF}}$ with these uncertainties. The reweighting function for the synthetic data set in Eq. 31 has the same form for HF and LF variations and is constructed as a ratio
@@
For $k=\textrm{HF}$, the b and c-quark efficiencies are varied, while for $k=\textrm{LF}$, we vary the efficiencies of light quarks and the gluons. We use the values $\nu_{k}=\pm 1$ for constructing synthetic data sets and fit the linear models
@@
The resulting parametrization is shown in Fig. 8. The HF and LF uncertainties are similar in shape. The HF variations range within 2%–5%, and the LF variations within 4%–8%. This slightly higher impact of the LF variations can be understood from the higher light-jet multiplicity after the $m(\mathrm{t}\overline{\mathrm{t}})\geq 750~{}\,\text{GeV}$ requirement.
@@ 8.5 Uncertainties in lepton efficiencies
Uncertainties in lepton efficiencies are described in Refs. [80, 81, 82] and handled by the weighting function in Eq. 34. Because the efficiency scale factors and corresponding uncertainties depend on the candidate’s pseudo-rapidity, which is not part of $\boldsymbol{x}$, we keep the $\boldsymbol{z}$ dependence on the l.h.s. and have
@@
which we use to define two surrogate data sets corresponding to the ±$\pm 1\sigma$ variations. Technically, we have $\mathcal{V}=\{-1,1\}$ and we learn a surrogate
@@
The variations are smaller than 1% in all cases, and the $\boldsymbol{x}$-dependence is almost negligible as is shown in Fig. 9.
@@ 8.6 Testing the tree-based estimates with neural networks
The comparisons in the preceding sections are given in terms of one-dimensional projections. A more general “Classifier two-sample test” (C2ST) [83, 84] can be used to check whether the parametric regression tree is fully expressive in the high-dimensional $\boldsymbol{x}$ space.
@@
The C2ST is a non-parametric method for determining whether two samples originate from the same distribution. The procedure involves training a binary classifier on a dataset comprising the two samples, with labels indicating their origin. The classifier’s performance is given by its accuracy and is used to infer the similarity between the distributions of the two samples. A better-than-chance performance suggests that the samples are drawn from different distributions. We take a specific nuisance parameter ${\boldsymbol{\nu}}$ and a pair of synthetic data sets, $\mathcal{D}_{\textrm{SM}}$ and $\mathcal{D}_{\nu}$ with $\nu\neq 0$. If a candidate estimate $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})$ is perfectly trained and fully expressive, we have
$\hat{S}(\boldsymbol{x}|\nu)=\frac{{\textrm{d}}\sigma({\boldsymbol{\nu}})\;\;\;\,}{{\textrm{d}}\sigma(\textrm{SM})}$ (118)
for all $\boldsymbol{x}$ and ${\boldsymbol{\nu}}$. It then follows that a new data set, defined in terms of $\mathcal{D}_{\boldsymbol{\nu}}$ by
@@
should be indistinguishable from $\mathcal{D}_{\textrm{SM}}$; because we reweight by the inverse of the presumably accurate estimate of the DCR, $\mathcal{D}_{\nu}$ should look like $\mathcal{D}_{\textrm{SM}}$. The idea is to test whether a classifier can distinguish $\mathcal{D}_{\textrm{reweighted}}$ from $\mathcal{D}_{\textrm{SM}}$ by evaluating its accuracy and attributing a p-value via constructing the distribution of the accuracy under the hypothesis that the samples are drawn from the same distribution.
We first train a classifier to distinguish the $\mathcal{D}_{\textrm{reweighted}}$ and $\mathcal{D}_{\textrm{SM}}$. We choose the nuisance related to HF b-tagging, $\nu=\nu_{\textrm{HF}}=1$ and thus test $\hat{S}_{\textrm{HF}}(\boldsymbol{x}|\nu_{\textrm{HF}}=1)$. The discriminator’s architecture is chosen differently than the surrogate’s. We take a neural network, implemented in pytorch, with sigmoid activation and three hidden layers of sizes 512, 512, and 256 and train it on half the data sets with the Adam optimizer. The resulting accuracy is 0.5001, suggesting almost perfect agreement. To assess the sensitivity of the test, we obtain the distribution of the accuracy under the hypothesis that the two samples are identical by merging $\mathcal{D}_{\textrm{reweighted}}$ and $\mathcal{D}_{\textrm{SM}}$ and randomizing the labels. We construct 1000 randomized pairs of identically distributed subsets from this randomized sample and train a classifier for each pair. The distribution of the accuracies is narrowly peaked and centered at 0.5, as shown in Fig. 10. Finally, we train another classifier to distinguish between the nominal sample $\mathcal{D}_{\nu_{\textrm{HF}}=1}$ and $\mathcal{D}_{\textrm{SM}}$. This time, the accuracy is 0.504, a highly significant deviation as shown in Fig. 10. The deviation from 0.5 is nevertheless numerically small, a result of the small $\boldsymbol{x}$-dependence of the DCR. Nevertheless, the neural-network classifier can distinguish ${\boldsymbol{\nu}}_{\textrm{HF}}=0$ and ${\boldsymbol{\nu}}_{\textrm{HF}}=1$ with highly significant accuracy. In summary, if we use our tree-based surrogate to remove $\nu_{\textrm{HF}}$-dependence from the simulation, a highly sensitive neural network is unable to distinguish the resulting distribution from the nominal SM simulation, suggesting excellent performance in the whole feature space.
@@ 8.7 Expected Limits from unbinned Asimov data
The Asimov data set [85] is frequently used to obtain expected exclusion limits from binned Poisson likelihoods [51]. The generalization to the unbinned case, discussed in Ref. [8], allows obtaining sampling-free exclusions in continuous parametric models. Our hypotheses are composite, defined by non-zero values of two Wilson coefficients. In the context of exclusion limits, we denote the null hypothesis by ${\boldsymbol{\theta}}$ with $N_{{{\boldsymbol{\theta}}}}=2$. The rest of the Wilson coefficients are profiled, i.e., treated as floating nuisance parameters. The alternate hypothesis corresponds to ${{\boldsymbol{\theta}}}={\mathbf{0}}$, again with $N_{{{\boldsymbol{\theta}}}}=2$.
Wilk’s theorem states that if the data are distributed under the (null) hypothesis $\boldsymbol{\theta}$, then the distribution of the test statistic $p(q_{{\boldsymbol{\theta}}}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ asymptotically agrees with a central $\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}$ distribution with $N_{{{\boldsymbol{\theta}}}}$ degrees of freedom. This distribution, therefore, is independent of the true value of the nuisance parameters. Because $q_{{\boldsymbol{\theta}}}$ is monotonic with the p-value, we can use it to define acceptance regions for ${\boldsymbol{\theta}}$ with sizes $\alpha=32\%$ or $\alpha=5\%$, corresponding to a CL of 68% and 95%, respectively. We expect to exclude a hypothesis ${\boldsymbol{\theta}}$ at a specific CL if there is at least a 50% chance for $q_{{\boldsymbol{\theta}}}$ to lie outside the corresponding acceptance region in case the (alternate) hypothesis ${{\boldsymbol{\theta}}}={\mathbf{0}}$ is true. Therefore, we must solve
@@
The final ingredient is Wald’s theorem [86], from which the distribution $\textrm{p}(q_{{\boldsymbol{\theta}}}|{\mathbf{0}})$ asymptotically follows as a non-central $\chi^{2}$-distribution with $N_{{\boldsymbol{\theta}}}$ degrees of freedom and non-centrality parameter $\Lambda$. This parameter can be computed (see Ref. [8] for details) for the unbinned likelihood ratio. In the same way, as for the binned case, it amounts to the Asimov expectation of Eq. 11 for the alternate hypothesis multiplied by -2. In our notation, the result is
@@
with the luminosity from Eq. 108 and the model’s DCR $\mathcal{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ given by Eq. 109. The sum extends over all events in the nominal $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ sample passing the event selection. The ratio $\mathcal{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ appears in the logarithm and the “extended” term corresponding to the total fiducial cross-section. We can obtain the test statistic under the alternate hypothesis and determine the expected exclusion contour from the profiled likelihood test statistic with this expression. The minimization is performed with the iminuit package [87].

### v3

@@ 7.2 Parton-level uncertainties
Figure 5: Variations of the simulation of the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ simulation with the nuisance parameters $\nu_{R}$ and $\nu_{F}$, modifying the renormalization and factorization scales, respectively. The dashed lines show the training data, and the solid lines show the prediction from $\hat{S}_{\textrm{scale}}(\boldsymbol{x}|\nu_{R},\nu_{F})$ as obtained from the BPT. All shapes are normalized to the simulation at the SM parameter point, corresponding to $\nu_{R}=\nu_{F}=0$.
Among the largest systematic effects are uncertainties in the factorization and renormalization scales, as detailed in Sec. 4.2. Following Eq. 16, setting $\nu_{R}=\pm 1$ and $\nu_{F}=\pm 1$ varies the scales $\mu_{R}$ and $\mu_{F}$ by a factor of 2. From simulation, we obtain event weights for all eight scale combinations,
@@
with the nominal SM simulation at $\nu_{R}=\nu_{F}=0$. We model the scale uncertainties up to quadratic accuracy in the nuisance parameters with the ansatz
@@
The five independent terms, labeled by $A={R,F,RR,FF,RF}$, cover the two linear, two quadratic, and one mixed term. The eight variations in $\mathcal{V}$ thus overconstrain these five functions.
We fit the BPT from Sec. 5.4 in the standard configuration from Sec. 7.1. The result is shown in Fig. 5 with one-dimensional projections for the observables most sensitive to variations in $\nu_{R}$ and $\nu_{F}$. Correlated scale variations by a factor of 2 for $\mu_{R}$ and $\mu_{F}$, corresponding to $\nu_{R}=\nu_{F}=\pm 1$ reach 8-10%. The exception is $p_{\textrm{T}}$, where the tail shows variations exceeding 20%. In this range, the fit has a small deficit of 1-2% relative to true variations, likely due to residual inflexibility in the quadratic model. Since $\mu_{R}$ and $\mu_{F}$ capture uncertainties from limited perturbative control, we ignore this slight mismatch for now. Other kinematic features show less shape dependence.
@@
We also simulate events at the SM parameter point using the alternative powheg generator, producing a similarly sized event sample for the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ process at NLO accuracy in the strong coupling constant. As outlined in Sec. 5.3, and with the caveats noted there, we assign a nuisance parameter $\nu_{\textrm{POW}}$, with $\nu_{\textrm{POW}}=0$ representing MadGraph5_amc@nlo and $\nu_{\textrm{POW}}=1$ representing powheg. Here, $\mathcal{V}=\{1\}$ allows us to train a single-parameter linear surrogate for the (log-) DCR, denoted $\hat{S}_{\textrm{POW}}(\boldsymbol{x}|\nu_{\textrm{POW}})$. Figure 6 shows one-dimensional projections of the features, revealing shape differences. Minor statistical fluctuations appear in the data tails due to the stochastic independence of the samples, but the BPT averages them out.
Uncertainties in the PDFs, which would require around 100 nuisance parameters for variations along the PDF eigendirections [79], are deferred for future treatment. Instead, and to account for uncertainties in the $m(\mathrm{t}\overline{\mathrm{t}})$ selection efficiency, we include a normalization uncertainty of 15%, setting $\alpha_{\textrm{norm}}=1.15$.
@@ 7.3 Jet energy calibration uncertainties
To evaluate the impact of uncertainties in the reconstructed transverse jet momenta, we vary the Delphes-predicted values according to the “total” CMS JES uncertainty provided in Ref. [62]. These variations affect the event selection, missing energy, top quark kinematic reconstruction, and other event features in $\boldsymbol{x}$. Since the per-jet variations depend on the nominal $p_{\textrm{T}}$ and pseudo-rapidity, which are latent (not in $\boldsymbol{x}$), the resulting function $J_{\nu_{\textrm{JES}}}(\boldsymbol{x},\boldsymbol{z})$ is also dependent on the latent event configuration.
Using the method in Sec. 4.4.1, we define synthetic data sets and set $\mathcal{V}=\{-1,-0.5,0.5,1\}$ for the JES nuisance parameter $\nu_{\textrm{JES}}$, parameterizing the per-jet variation effects on $\boldsymbol{x}$ in units of the JES uncertainty standard deviations. The half-integer values for $\nu_{\textrm{JES}}$ provide more granularity than the typical ±$\pm 1\sigma$ variations used in binned LHC analyses. We then fit a log-linear surrogate,
@@
to model the JES dependence. Figure 7 shows an excellent fit of the surrogate to the variations in the training data. Most observables show a flat variation, except for $p_{\textrm{T}}(\mathrm{t}\overline{\mathrm{t}})$, which rises from 2%–5%. In the tails of $m(\mathrm{t}\overline{\mathrm{t}})$, slight asymmetries in the training data variations are not captured by the linear model, as it approximately symmetrizes the total uncertainty. Refinement with a higher-degree surrogate is left for future work.
@@ 7.4 Uncertainties in tagging efficiencies
Uncertainties in the b-tagging efficiency for jets, along with their application, are provided in Ref. [62]. This approach relies on $p_{\textrm{T}}$, pseudo-rapidity, and a nominal binary b-tag label from Delphes. To apply variations, we also need the generator-level jet flavor $f$ within $\{\textrm{udsg},\mathrm{c},\mathrm{b}\}$. Using the nominal Delphes simulation, we parametrize the $p_{\textrm{T}}$ and $\eta$-dependent $\mathrm{b}$-tagging efficiencies $\varepsilon_{f}(p_{\textrm{T}},\eta)$ for each flavor. Two systematic uncertainties are considered with scale factors $\textrm{SF}_{f}(p_{\textrm{T}},\eta)$ and variations $\Delta\textrm{SF}_{f}(p_{\textrm{T}},\eta)$. The heavy-flavor (HF) tagging uncertainty covers the b and c-quark tagging rates, modified in a correlated way for non-zero $\nu_{\textrm{HF}}$. The light-flavor (LF) mistagging uncertainty addresses tagging rates for light-quark and gluon jets, associated with $\nu_{\textrm{LF}}$. The reweighting function for synthetic data in Eq. 31 is given by
@@
For $k=\textrm{HF}$, we vary b and c-jet efficiencies, and for $k=\textrm{LF}$, we vary the efficiencies for light-quark and gluon jets. Using $\nu_{k}=\pm 1$, we construct synthetic data sets and fit linear surrogates
@@
The resulting parametrization is shown in Fig. 8. The HF and LF uncertainties exhibit similar shapes. HF variations range from 2% to 5%, while LF variations show a slightly larger impact, ranging from 4% to 8%. This greater effect of the LF variations is due to the higher light-jet multiplicity following the $m(\mathrm{t}\overline{\mathrm{t}})\geq 750~{}\,\text{GeV}$ selection.
@@ 7.5 Uncertainties in lepton efficiencies
Uncertainties in lepton efficiencies are detailed in Refs. [80, 81, 82] and are handled using the weighting function in Eq. 34. Since the efficiency scale factors and uncertainties depend on the candidate’s pseudo-rapidity, which is not included in $\boldsymbol{x}$, we retain the $\boldsymbol{z}$ dependence in
@@
used to define two surrogate data sets corresponding to ±$\pm 1\sigma$ variations. Here, $\mathcal{V}=\{-1,1\}$, and we learn a surrogate
@@
Variations are under 1% in all cases, with minimal $\boldsymbol{x}$-dependence, as shown in Fig. 9.
@@ 7.6 Testing the tree-based estimates with neural networks
The comparisons in previous sections are one-dimensional projections. For a more general check of whether the BPT is fully expressive in the high-dimensional $\boldsymbol{x}$ space, we can apply a “Classifier two-sample test” (C2ST) [83, 84].
@@
The C2ST is a non-parametric method for assessing if two samples originate from the same distribution. It trains a binary classifier on a combined dataset of the two samples, using labels to indicate sample origin. The classifier’s accuracy reveals distribution similarity; accuracy above chance suggests different distributions.
Given a specific nuisance parameter ${\boldsymbol{\nu}}$ and a pair of synthetic data sets, $\mathcal{D}_{\textrm{SM}}$ and $\mathcal{D}_{\nu}$ with $\nu\neq 0$, if a candidate estimate $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})$ is accurate and fully expressive, then
$\hat{S}(\boldsymbol{x}|\nu)=\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})\;\;\,\,}{{\textrm{d}}\sigma(\boldsymbol{x}|\textrm{SM})}$ (91)
for all $\boldsymbol{x}$ and ${\boldsymbol{\nu}}$. Thus, reweighting $\mathcal{D}_{\boldsymbol{\nu}}$ to form
@@
should make $\mathcal{D}_{\textrm{reweighted}}$ indistinguishable from $\mathcal{D}_{\textrm{SM}}$. To test this, a classifier’s accuracy in distinguishing $\mathcal{D}_{\textrm{reweighted}}$ from $\mathcal{D}_{\textrm{SM}}$ is used, with a p-value based on the null distribution of the accuracy. We train a classifier using HF b-tagging with $\nu=\nu_{\textrm{HF}}=1$ to test $\hat{S}_{\textrm{HF}}(\boldsymbol{x}|\nu_{\textrm{HF}}=1)$. The classifier, a neural network in pytorch with sigmoid activation and three hidden layers (512, 512, 256 units), is optimized with Adam on half of the data. Its accuracy is 0.5001, suggesting near-perfect agreement. To evaluate this result, we merge $\mathcal{D}_{\textrm{reweighted}}$ and $\mathcal{D}_{\textrm{SM}}$, randomize labels, and train 1000 classifiers on pairs of identical subsets. The null distribution peaks at 0.5, as shown in Fig. 10. For comparison, distinguishing $\mathcal{D}_{\nu_{\textrm{HF}}=1}$ from $\mathcal{D}_{\textrm{SM}}$ yields 0.504, a significant deviation (Fig. 10). This deviation is small due to the mild $\boldsymbol{x}$-dependence of the DCR, yet the neural network accurately detects the difference. In summary, removing $\nu_{\textrm{HF}}$-dependence with our surrogate makes it impossible for a high-sensitivity neural network to distinguish from the SM, indicating strong performance across the feature space.
@@ 7.7 Expected Limits from unbinned Asimov data
The Asimov dataset [85] is commonly used to derive expected exclusion limits from binned Poisson likelihoods [59]. Ref. [8] extends this to the unbinned case, enabling sampling-free exclusions within continuous parametric models. Here, we consider composite hypotheses involving two Wilson coefficients, which we denote by ${\boldsymbol{\theta}}$. Under the exclusion scenario, ${\boldsymbol{\theta}}$ represents the null hypothesis with $N_{{{\boldsymbol{\theta}}}}=2$, while other Wilson coefficients are profiled as nuisance parameters. The alternative hypothesis assumes ${{\boldsymbol{\theta}}}={\mathbf{0}}$.
Wilks’ theorem states that if the data are distributed under the null hypothesis $\boldsymbol{\theta}$, the test statistic $p(q_{{\boldsymbol{\theta}}}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ asymptotically follows a central $\chi^{2}$ distribution with $N_{{{\boldsymbol{\theta}}}}$ degrees of freedom. This distribution is independent of the true values of the nuisance parameters. Given that our POIs primarily influence the predictions linearly [74], we assume any minor quadratic terms do not invalidate Wilks’ theorem [27]. However, in practical applications, this assumption should be verified, as shown in Ref. [8], where good agreement was observed. Since $q_{{\boldsymbol{\theta}}}$ is monotonic with the p-value, it can define acceptance regions for ${\boldsymbol{\theta}}$ at confidence levels of 68% ($\alpha=32\%$) or 95% ($\alpha=5\%$). We anticipate excluding a hypothesis ${\boldsymbol{\theta}}$ at a given CL if there’s a 50% or greater probability for $q_{{\boldsymbol{\theta}}}$ to fall outside the corresponding acceptance region when the alternate hypothesis ${{\boldsymbol{\theta}}}={\mathbf{0}}$ is true. Therefore, we must solve
@@
The final ingredient is Wald’s theorem [86], which implies that the distribution $\textrm{p}(q_{{\boldsymbol{\theta}}}|{\mathbf{0}})$ asymptotically follows a non-central $\chi^{2}$ distribution with $N_{{\boldsymbol{\theta}}}$ degrees of freedom and non-centrality parameter $\Lambda$. This parameter can be computed (see Ref. [8] for details) for the unbinned likelihood ratio. As in the binned case, it corresponds to the Asimov expectation of Eq. 11 for the alternate hypothesis, multiplied by $-2$. In our notation, the result is
@@
with the integrated luminosity from Eq. 81 and the model’s DCR $\mathcal{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ from Eq. 82. The sum is over all events in the nominal $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ sample passing the event selection. The ratio $\mathcal{R}(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ appears in both the logarithm and the “extended” term for the total fiducial cross-section. This expression provides the test statistic under the alternate hypothesis, allowing us to obtain the expected exclusion contour from the profiled likelihood test statistic. The minimization is performed with the iminuit package [87].